Determine the soil water content in percentage by weight for a 120g sample in a moist condition that, when oven-dried, weighs 95g

What additional information do you need to determine porosity?
Answer #1

Here given

Weight of soil in moist condition is = 120 g.

After oven dried soil weight is = 95 g.

So water present in soil = 120 - 95 = 25 g.

( Moist weight - dry weight)

In terms of percentage = (25 ÷ 120) × 100 = .20833 × 100 = 20.83%

To determine porosity , volume of total soil and particle density of soil require.

From total volume and total mass of soil we can calculate bulk density of soil.

Bulk density = total wright of soil ÷ total volume.

So porosity is = (particles density - bulk density) ÷ particles density.